What is the probability that two female students will be selected at random to
participate in a certain research project, from a class of seven males and three
female students?
1
Expert's answer
2020-11-22T17:03:34-0500

Total number of student in a class"=7+3=10"

Number of ways of selecting two student from 10 student "= \ ^{10}C_2=\frac{10!}{2.(10-2)!}=\frac{10.9}{2}=45"

Number of ways of selecting 2 girls from 3 girls"= \ ^3C_2=\frac{3!}{2.1!}=3"

The probability that both student are girls "=\frac{3}{45}=\frac{1}{15}"